# Auto-Landmark: Acoustic Landmark Dataset and Open-Source Toolkit for Landmark Extraction thanks: Our Dataset Can be found in https://github.com/Tonyyouyou/Landmark_Dataset

### v1

@@ Abstract
I Introduction
In speech processing, frame-based methods are most commonly used to segment the speech waveform. This method treats each frame as the central unit of analysis, with a fixed set of speech attributes measured at each frame. However, because this approach relies on fixed-duration processing, it often overlooks important timing factors such as speaking rate and segmental duration. In contrast, acoustic landmark detection focuses on specific, acoustically significant points in the speech signal [1], as illustrated in Figure 1. This method allows for the identification of acoustic landmarks somewhat independently of frames, while also providing valuable timing information for subsequent processing.
@@ I Introduction
II Related Work
II-A Acoustic Landmarks
@@
II-B Application of Acoustic Landmarks
TABLE I: Description of the five landmarks investigated.
@@ II-B Application of Acoustic Landmarks
Acoustic landmarks were initially used in speech recognition [1, 2], with some researchers exploring the combination of acoustic landmarks with CTC in this context [2]. Following their success in speech recognition, acoustic landmarks have been applied to various other domains. Acoustic landmarks have been widely applied in various health-related studies. For example, [3, 4] utilized acoustic landmark timing information for depression detection. [7] further extended the use of acoustic landmarks by tokenizing speech for large language model-based systems in depression detection. [6] used acoustic landmark time duration to describe abnormalities in speech production that affect a speaker’s intelligibility. [5] evaluated the feasibility of differentiating conversational and clear speech produced by individuals with muscle tension dysphonia (MTD) using landmark-based analysis of speech.
III Acoustic Landmark Dataset
III-A Dataset and Landmark Selection
As shown in Table I, we selected five different types of acoustic landmarks based on previous research [7, 3, 8, 4, 5, 6], each with the onset and offset states. These landmarks are g(lottis), s(sonorant), f(ricative), v(voiced fricative), and b(ursts), which are pointing in time where different abrupt articulatory events occur.
@@ III-B Labeling Convention
TABLE II: Rules for Annotating Data
@@
Table II outlines the rules we followed for annotating acoustic landmarks. Each landmark corresponds to specific phonetic events and transitions within the speech signal. The v+ and v- landmarks indicate the start and end times of voiced fricatives, such as [z], [zh], [v], and [dh]. Similarly, f+ and f- mark the start and end times of fricatives like [s], [sh], [f], and [th]. For nasals and the glides [l], the s+ and s- landmarks represent the start and end times, covering phonemes like [m], [n], [ng], [em], [en], [eng], and [nx]. The b+ and b- landmarks capture the sudden energy changes in stops and affricates, including [b], [d], [g], [p], [t], [k], [dx], [q], [jh], and [ch] and manually labeled. Finally, the g+ and g- landmarks denote the start and end times of voiced phones, encompassing a range of vowels (e.g., [iy], [ih], [eh], [ey], [ae], [aa], [aw], [ay], [ah], [ao], [oy], [ow], [uh], [uw], [ux], [er], [ax], [ix], [axr], [ax-h]), voiced fricatives, nasals, and stops such as [b], [d], and [g]. The g- landmark also indicates the transition from voiced to voiceless sounds.
III-C Statistics of Landmark Dataset
@@
IV Landmark Detection System
@@
IV-A Filter-based Method
@@
IV-B Deep Learning Based Method
To date, landmark detection has not been performed with deep learning models. Our first consideration is how to model this problem. In this paper, since both landmark detection and speech recognition tasks involve converting speech signals into symbols, we approach landmark detection as a speech recognition task. We employ a Hybrid CTC/attention end-to-end architecture [16]. We utilized two types of models as the Encoder: the Conformer [17] and ConExtBimamba [18]. For the decoder, we employed the Transformer model [19]. Given the success of self-supervised models in various speech tasks, we therefore explored how they perform on the landmark detection task. Specifically, we utilized the wav2vec2.0 large model [20]. We followed the Superb setup [21], using the weighted sum feature from different layers of wav2vec2.0.
@@ IV-B Deep Learning Based Method
IV-C Toolkit for Landmark Extraction
@@
V Results and Discussion
V-A Experimental Setup
@@
TABLE III: Landmark detection results for different methods
@@
V-B Baseline Result
Table III compares the landmark detection results from the two methods, evaluated by Landmark Error Rate (LER). LER was calculated based on patterns without considering landmark timing. Additionally, following previous conventions, we did not differentiate between onset (+) and offset (–) landmarks [1].All the deep learning models produced lower LER compared to both signal processing methods, with the ConExtBiMamba model outperforming the others. Typically, the overall performance of models improves when using self-supervised models. However, as shown in Table III, the performance of both the Conformer and ConExtBimamba models decreased when combined with wav2vec2, contrary to expectations. This issue arose because several landmarks occur in close temporal proximity to each other, even simultaneously. This means that numerous landmarks can appear within a 10ms window. However, as wav2vec uses a 20ms window with a 10ms stride, it potential misses some landmarks.
V-C Insights from Self-supervised Weights

### v2

@@ Abstract
1 Introduction
In speech processing, frame-based methods are most commonly used to segment the speech waveform. This method treats each frame as the central unit of analysis, with a fixed set of speech attributes measured at each frame, has been used in many domain [1, 2]. However, because this approach relies on fixed-duration processing, it often overlooks important timing factors such as speaking rate and segmental duration. In contrast, acoustic landmark detection focuses on specific, acoustically significant points in the speech signal [3], as illustrated in Figure 1. This method allows for the identification of acoustic landmarks somewhat independently of frames, while also providing valuable timing information for subsequent processing.
@@ 1 Introduction
2 Related Work
2.1 Acoustic Landmarks
@@
2.2 Application of Acoustic Landmarks
Table 1: Description of the five landmarks investigated.
@@ 2.2 Application of Acoustic Landmarks
Acoustic landmarks were initially used in speech recognition [3, 4], with some researchers exploring the combination of acoustic landmarks with CTC in this context [4]. Following their success in speech recognition, acoustic landmarks have been applied to various other domains. Acoustic landmarks have been widely applied in various health-related studies. For example, [5, 6] utilized acoustic landmark timing information for depression detection. [9, 14, 15] further extended the use of acoustic landmarks by tokenizing speech for large language model-based systems in depression detection. [8] used acoustic landmark time duration to describe abnormalities in speech production that affect a speaker’s intelligibility. [7] evaluated the feasibility of differentiating conversational and clear speech produced by individuals with muscle tension dysphonia (MTD) using landmark-based analysis of speech.
3 Acoustic Landmark Dataset
3.1 Dataset and Landmark Selection
As shown in Table 1, we selected five different types of acoustic landmarks based on previous research [9, 5, 10, 6, 7, 8], each with the onset and offset states. These landmarks are g(lottis), s(sonorant), f(ricative), v(voiced fricative), and b(ursts), which are pointing in time where different abrupt articulatory events occur.
@@ 3.2 Labeling Convention
Table 2: Rules for Annotating Data
@@
Table 2 outlines the rules we followed for annotating acoustic landmarks. Each landmark corresponds to specific phonetic events and transitions within the speech signal. The v+ and v- landmarks indicate the start and end times of voiced fricatives, such as [z], [zh], [v], and [dh]. Similarly, f+ and f- mark the start and end times of fricatives like [s], [sh], [f], and [th]. For nasals and the glides [l], the s+ and s- landmarks represent the start and end times, covering phonemes like [m], [n], [ng], [em], [en], [eng], and [nx]. The b+ and b- landmarks capture the sudden energy changes in stops and affricates, including [b], [d], [g], [p], [t], [k], [dx], [q], [jh], and [ch] and manually labeled. Finally, the g+ and g- landmarks denote the start and end times of voiced phones, encompassing a range of vowels (e.g., [iy], [ih], [eh], [ey], [ae], [aa], [aw], [ay], [ah], [ao], [oy], [ow], [uh], [uw], [ux], [er], [ax], [ix], [axr], [ax-h]), voiced fricatives, nasals, and stops such as [b], [d], and [g]. The g- landmark also indicates the transition from voiced to voiceless sounds.
3.3 Statistics of Landmark Dataset
@@
4 Landmark Detection System
@@
4.1 Filter-based Method
@@
4.2 Deep Learning Based Method
To date, landmark detection has not been performed with deep learning models. Our first consideration is how to model this problem. In this paper, since both landmark detection and speech recognition tasks involve converting speech signals into symbols, we approach landmark detection as a speech recognition task. We employ a Hybrid CTC/attention end-to-end architecture [20]. We utilized two types of models as the Encoder: the Conformer [21] and ConExtBimamba [22, 23, 24]. For the decoder, we employed the Transformer model [25]. Given the success of self-supervised models in various speech tasks, we therefore explored how they perform on the landmark detection task. Specifically, we utilized the wav2vec2.0 large model [26]. We followed the Superb setup [27], using the weighted sum feature from different layers of wav2vec2.0.
@@ 4.2 Deep Learning Based Method
4.3 Toolkit for Landmark Extraction
@@
5 Results and Discussion
5.1 Experimental Setup
@@
Table 3: Landmark detection results for different methods
@@
5.2 Baseline Result
Table 3 compares the landmark detection results from the two methods, evaluated by Landmark Error Rate (LER). LER was calculated based on patterns without considering landmark timing. Additionally, following previous conventions, we did not differentiate between onset (+) and offset (–) landmarks [3].All the deep learning models produced lower LER compared to both signal processing methods, with the ConExtBiMamba model outperforming the others. Typically, the overall performance of models improves when using self-supervised models. However, as shown in Table 3, the performance of both the Conformer and ConExtBimamba models decreased when combined with wav2vec2, contrary to expectations. This issue arose because several landmarks occur in close temporal proximity to each other, even simultaneously. This means that numerous landmarks can appear within a 10ms window. However, as wav2vec uses a 20ms window with a 10ms stride, it potential misses some landmarks.
5.3 Insights from Self-supervised Weights